# Combining Static and Dynamic Approaches for Mining and Testing Constraints for RESTful API Testing

### v1

@@ 3.1. Constraints from Request Specifications
An essential component of an API specification is the request specification. This part instructs clients on how to correctly initiate API calls, detailing the required inputs and the returned data. Our constraint mining relies on natural language descriptions attached to request parameters or response properties. Since descriptions are optional in OAS, they may appear in one endpoint but not others. To extract them, we first check the current response schema; if none are found, we search the entire OAS document. Unlike KAT (Le et al., 2024), which uses LLM-based ”description mapping,” we avoid potential inaccuracies that could propagate errors. Instead, we adopt descriptions from properties with the same name in other schemas. If none exist, we exclude the property, prioritizing precision over recall.
3.1.1. Request-Response Constraint Mapping
The idea is that to determine a constraint from request parameters, the response data must contain a property that reflects this constraint. Thus, we identify pairs consisting of a request parameter that includes a constraint and a response data property that reflects this constraint (Figure 4). For instance, to verify a ”customerID” constraint from Observation 1, the response data should contain a property that identifies the customer. Thus, the required pair is ¡”customerID”, ”customer”¿. If the response data does not have a field to represent a constraint, that constraint is disregarded as it cannot be validated.
@@ 3.1.1. Request-Response Constraint Mapping
10If a matching property exists, explain it using this format: ...’’’
@@ 3.1.2. Detailed Process
Algorithm 1 Extract Constraints from Request Parameters
@@
Initially, the description extraction phase follows the process in Algorithm 1. An API request comprises multiple parameters and a response data schema. First, a prompt is made to gather observations on the response schema and operations associated with this request (lines 3–4, Algorithm 1). For a request parameter, we engage the LLM to provide its observations on that parameter by presenting it alongside its description (line 12, Algorithm 1). The LLM is expected to provide a description of a constraint within this parameter description. If the description is missing, the specification is searched for another parameter with an identical name.
These descriptions support the next step: Request-Response parameter mapping (Block 4, line 13). We guide the LLM through a two-step reasoning process using a tailored prompt (Figure 4). The LLM receives a brief description of the request parameter, including its details and observations ( 2), ensuring a clear understanding of its intent and constraints. Second, we present the response schema, observations from Block 3, and ask the LLM to identify a matching property. A match occurs when the request parameter filters response data or both share the same value meaning.
From this two-step reasoning, we require the LLM to answer three questions: (1) Is there a matching property in the response schema? (2) What is this property? (3) How does the request parameter influence this property?
If the answer to (1) is false, indicating that no property reflects this request parameter, we disregard this parameter. Otherwise, we proceed to the final prompt (Figure 5), which is the confirmation of the mapping (lines 14–17, Block 5). In this prompt, we present the pair of the request parameter and the matched property from Block 4 and ask the LLM to confirm the accuracy of this mapping (line 15). This step aims to minimize the occurrence of false positives. Finally, the validated pairs of the request parameters and the response properties are stored as Request-Response constraints (lines 16–17).
@@ 3.2. Constraints from Response Specifications
3:   for each prop in response_schema do
4:    desc $\leftarrow$ response_schema[prop][”description”]
5:    if desc = NULL then
6:     desc $\leftarrow$ exactMatchProp(API_spec, prop)
@@
8:      continue # Skip this property
9:    if prop in knowledge_base then
10:     if knowledge_base[prop] = TRUE then
11:      constraint_properties.append(prop)
12:    else
13:     datatype $\leftarrow$ response_schema[prop][”datatype”]
14:     prop_obser $\leftarrow$ LLM().propertyObser(prop, datatype, desc)
15:     constraint_confirmation $\leftarrow$ LLM().constraint_confirmation (prop, datatype, desc, prop_obser)
16:     if constraint_confirmation = TRUE then
17:      constraint_properties.append(prop)
18:     # Add this property to knowledge base
19:     knowledge_base[prop] $\leftarrow$ constraint_confirmation
20:   return constraint_properties
21:  end function
We examine descriptions within the response schema specification, as they provide direct constraints by mapping each property. Each endpoint includes a response schema that guides clients on parsing returned data, structuring the response object with properties and descriptions (Figure 2). Our constraint-mining algorithm, leveraging LLM, is outlined in Algorithm 2. It first extracts descriptions for each property (lines 4-8) following Section 3.1. If a description exists, we check a knowledge base of LLM-identified constraints to avoid redundant queries. If found, we reuse the stored information; otherwise, we prompt LLM to extract constraints.
Observation-Confirmation Strategy. This process involves two phases: observation (line 14) and confirmation (line 15). Our experiments reveal that when descriptions lack detail for constraint extraction, LLM may resort to fabricating details, a phenomenon known as hallucination. Drawing inspiration from the Chain-of-thought (Wei et al., 2022), we divide the task of extracting constraints into two phases of observation and confirmation, the initial prompt better contextualizes the description of constraints, enabling the subsequent prompt to more accurately determine a constraint.
In the observation phase (Block 6), LLM is prompted to identify constraints from the description. For example, given a property date (type: string) described as “ISO date: the literal date of the holiday”, LLM might infer: ”The date must follow the YYYY-MM-DD format for validity, ensuring consistency within the API.” This adds specificity not explicitly mentioned in the description.
@@
4. Combining Constraints and Invariants
Constraints detected by our static method are based on OAS while invariants determined by AGORA come from execution data. We present an ensemble approach to combine constraints and invariants. Figure 6 shows sample outputs produced by each method. Several constraints can only be identified using OAS. For instance, Constraint 8 in Figure 6, which specifies that the charge amount must be smaller than 99,999,999, can only be extracted from the OAS. This is because AGORA requires a sufficiently large number of API executions to infer such constraints. Conversely, some constraints can only be inferred at run-time, e.g., those related to returned Hrefs—URLs containing requested information.
Figure 6. Example outputs from RBCTest and AGORA.
@@ 4. Combining Constraints and Invariants
7
return.price must be
within input.price_range
@@ 8 An amount is a positive integer
In AGORA, each detected invariant is associated with a set of variables, as depicted in Figure 6, and a given set of variables may be linked to one or more invariants (e.g. Invariants 3-4). In our static method, each constraint is tied to a specific set of variables, which aligns with how AGORA groups invariants. Because each constraint is modeled via textual representations, we leverage an LLM and its natural-language text understanding to derive the relevant variables in the constraint.
If a constraint and an invariant involve different sets of variables, we include both in the resulting constraints for RBCTest, as they are uniquely detected by each approach. If they involve the same, we select the constraint or invariant with the stricter condition. For instance, in Figure 6, row 6, we chose the constraint identified by our static method, as it is stricter than the invariant detected by AGORA. To determine which is stricter, we leverage an LLM, which can interpret well the conditions expressed in natural language.
5. Constraint Test Generation
1CONSTRAINT_TEST_GEN_PROMPT = ’’
3Constraint description:
4- Constraint from request parameter: {parameter}
5- Constraint description: {constraint_description}
6API response schema: {response_schema_specification}
7The property of the provided request parameter in API response:
8- "{property}": "{prop_description}"
9Based on the provided constraint from request param, and the respective attribute in the API response, generate a Python script to verify the ’{property}’ property in the response.
10Rules: {Rules for test gen}
11Format the script as shown: ...’’’’
Figure 7. Constraint Test Generation Prompts
Constraint tests are generated using LLM (Figure 8) based on two types of constraints: Request-Response constraints and Response Property constraints. Request-Response constraints stem from request parameter descriptions, while Response Property constraints are derived from the response schema. These tests take a response body and request details as input, producing outcomes of ‘matched,’ ‘mismatched,’ or ‘unknown.’ A ‘matched’ outcome confirms that the response satisfies the constraint, whereas ‘mismatched’ indicates a violation. An ‘unknown’ outcome suggests the absence of the relevant property, possibly due to its optional nature in the specification. We also incorporate a semantic verifier that cross-checks test cases against examples in the OAS file. If a test case fails to validate a given correct example, it suggests that the mined constraint may be incorrect.
Figure 8. Constraint Test Generation in RBCTest
5.1. Request-Response Constraints Testing
The Request-Response Constraint Test identifies dependencies between a constrained request parameter and its corresponding response property. To generate a test case, we use the prompt in Figure 7, which requires four inputs: the parameter name, its constraint description, the corresponding property name, and the response schema. LLM generates a validation function with two inputs: the response body and the request parameter. It then creates a script to check conditions between them. For instance, when validating a ‘created’ time interval (Observation 1), LLM extracts the ‘created’ time from the response body and the conditional values from the request parameter (‘created[gte]’, ‘created[lte]’) before performing logical comparisons. To ensure robustness, we guide LLM with predefined rules: using a try-catch block for error handling, excluding examples, and following a standardized function template. These rules ensure consistency and focus on constraint verification.
5.2. Response Property Constraints Testing
The Response Property constraint Test directly correlates the property’s description with the property itself. To create a test case for this constraint, we utilize a specific prompt (Figure 7). This prompt requires three inputs: the property name, the constraint description, and the response data schema. The description extracted from the previous mining step gives the necessary information for generating constraint validation code, while the response data schema defines the structure and type of the expected data. This schema guides LLM in generating code to parse response data. We guide LLM to follow predefined rules (similar to 5.1) to maintain consistency in the generated code in different constraints.
6. Empirical Evaluation
For evaluation, we conduct several experiments, seeking to answer the following research questions:
RQ1. [Constraints Test Generation Accuracy] How well does RBCTest perform in comparison with individual static and dynamic approaches in generating test cases from the mined constraints?
RQ2. [Constraints Mining Accuracy] How well does RBCTest perform in detecting the onstraints for response bodies in the RESTful API specification?
RQ3. [Accuracy in Test Generation from the Correctly Mined Constraints] How accurate is our approach in generating test cases for mined constraints?
RQ4. [Usefulness in Response Body Testing] How accurate is our approach in detecting mismatches between the specification and its working APIs?
RQ5. [Ablation Study] How well do Confirmation-Observation and Semantic Verifier contribute to RBCTest’s performance?
Data Collection
We curated a dataset from eight real-world services, including GitLab and Stripe, comprising 59 endpoints and 83 operations. These services were selected for several reasons: (1) They feature complex request-response structures across diverse business domains. (2) They have been widely used in API testing research, such as ‘Canada Holidays’ (Le et al., 2024), GitLab-services (Yamamoto, 2021; Wu et al., 2022; Atlidakis et al., 2019; Karlsson et al., 2020; Lin et al., 2022; Le et al., 2024), and ‘Stripe’ (Le et al., 2024; Martin-Lopez et al., 2022; Mirabella et al., 2021). (3) Their active status allows API calls to collect real response data. (4) Their specifications vary in documentation quality, enabling evaluation across different levels of completeness. We selected API endpoints based on the presence of parameter or response descriptions, excluding those without any descriptions, as constraints could not be identified. Stripe, offering a test mode with limited endpoints, contains deeply nested response schemas, often missing values for validation. To mitigate this, we included only Stripe endpoints without nested schemas, retaining 7 in total. We cover 3 REST methods: GET, PUT, and POST (Table 3).
7. Experimental Results
7.1. Constraints Test Generation Accuracy (RQ1)
Table 1. Constraints Test Generation on AGORA dataset.
Static Dynamic Unique Overlapping
APIs Total TP P Total TP P Static Dynamic +S +D Eq.
A.Hotel 44 39 88.6 117 61 52.1 20 34 8 1 10
GitHub’ 16 13 81.3 198 194 98 3 175 6 0 4
GitHub” 7 7 100 150 127 84.7 7 121 0 0 0
Marvel 28 23 82.1 115 55 47.8 13 35 6 0 4
OMDB 4 4 100 16 15 93.8 3 13 0 0 1
OMDB’ 2 1 50 5 5 100 0 3 0 0 1
Spotify 21 21 100 41 41 100 11 26 3 2 5
Spotify’ 13 12 92.3 68 58 85.3 1 29 5 4 2
Spotify” 15 15 100 55 45 81.8 3 28 6 2 4
Yelp 2 1 50 30 12 40 1 11 0 0 0
YouTube 65 62 95.4 194 111 57.2 45 52 12 2 3
TOTAL 217 198 91.2 989 724 73.2 107 527 46 11 34
  • Static: LLM-based, Dynamic: Execution-based (AGORA), +S: Static constraint better, +D: Dynamic constraint better, Eq: Equivalent.

### v3

@@ 5.1. Constraints from Request Specifications
Figure 3 describes the process of mining constraints. Constraints for response bodies may exist in descriptions of operations, parameters, and other information defined in the request specifications.
@@
10If a matching property exists, explain it in this format:...’
@@
5.1.1. Request-Response Constraint Mapping
The response data may contain a property that reflects a constraint available in request parameters. Thus, we identify pairs consisting of a request parameter that includes a constraint and a response data property that reflects this constraint (Figure 4). For instance, to verify a ”customerID” constraint, the response data should contain a property that identifies the customer. Thus, the required pair is ¡”customerID”, ”customer”¿. If the response data does not have a field to represent a constraint, that constraint is disregarded as it cannot be validated.
@@ 5.1.2. Detailed Process
Algorithm 1 Extract Constraints from Request Specifications
@@
Algorithm 1 provides the details. For each API request, a prompt is made to gather observations on the response schema and operations associated with this request (lines 3–4, Algorithm 1). With each request parameter (line 5), we engage the LLM to offer observations on the parameter using its description (line 12, Algorithm 1).
These parameter observations support the next step: Request-Response parameter mapping (Block 4, line 13). We guide the LLM through a two-step reasoning process using a tailored prompt (Figure 4). The LLM receives a brief description of the request parameter, including its details and observations ( 2), ensuring an understanding of its intent and constraints. We then present the response schema, observations from Block 3, and ask the LLM to identify a matching property. A match occurs when the request parameter filters response data or both share the same meaning. From this two-step reasoning, we require the LLM to answer three questions: (1) Is there a matching property in the response schema? (2) What is this property? (3) How does the request parameter influence this property?
If the answer to (1) is false, i.e., no property reflects this request parameter, we disregard this parameter. Otherwise, we proceed to the final prompt (Figure 5), which is the confirmation of the mapping (lines 14–17, Block 5). We present the pair of the request parameter and the matched property from Block 4, and ask the LLM to confirm the accuracy of this mapping (line 15). This step aims to minimize the occurrence of false positives. Finally, the validated pairs of request parameters and response properties are stored as Request-Response constraints (lines 16–17).
@@ 5.2. Constraints from Response Specifications
3:     for each prop in resp_schema do
4:        desc $\leftarrow$ resp_schema[prop][”description”]
5:     +resp_schema[prop][”type”]+resp_schema[prop][”format”]
6:     +resp_schema[prop][”example”]
@@
8:           desc $\leftarrow$ exactMatchProp(API_spec, prop)
9:           if desc = NULL then
10:              continue # Skip this property
11:        if prop in knowledge_base then
12:           if knowledge_base[prop] = TRUE then
13:              constraint_properties.append(prop)
14:        else
15:           datatype $\leftarrow$ response_schema[prop][”datatype”]
16:           prop_obser $\leftarrow$ LLM().propertyObser(prop, datatype, desc)
17:           constraint_confirmation $\leftarrow$ LLM().constraint_confirmation (prop, datatype, desc, prop_obser)
18:           if constraint_confirmation = TRUE then
19:              constraint_properties.append(prop)
20:           # Add this property to knowledge base
21:           knowledge_base[prop] $\leftarrow$ constraint_confirmation
22:     return constraint_properties
23:  end function
We examine descriptions, data formats, and examples in the response schema specification, as they may provide constraints for a property. Each endpoint includes a response schema that guides clients on parsing returned data, structuring the response object with properties, descriptions, data formats, and examples (Figure 2).
Algorithm 2 shows our constraint-mining algorithm for response specifications. It first extracts descriptions, data formats, and examples for a property (lines 4-10). If such information exists, we check the knowledge base of LLM-identified constraints to avoid redundant queries (lines 11-13). If found, we reuse the stored data; otherwise, we prompt the LLM to extract constraints (lines 15-21).
Observation-Confirmation Strategy. This process involves two phases: observation (line 16) and confirmation (line 17). Our experiments reveal that when descriptions lack detail for constraint extraction, the LLM may resort to fabricating details, a phenomenon known as hallucination. Drawing inspiration from the Chain-of-thought (chain_of_thought), we divide the task of extracting constraints into two phases of observation and confirmation, the initial prompt better contextualizes the description of constraints, enabling the subsequent prompt to more accurately determine a constraint.
In the observation phase (Block 6), the LLM is prompted to identify constraints from the description. For example, given a property date (type: string) described as “ISO date: the literal date of the holiday”, the LLM might infer: ”The date must follow the YYYY-MM-DD format for validity, ensuring consistency within the API.”
@@
6. Constraint Test Generation
1CONSTRAINT_TEST_GEN_PROMPT = ’’
3Constraint description:
4- Constraint from request parameter: {parameter}
5- Constraint description: {constraint_description}
6API response schema: {response_schema_specification}
7The property of the request parameter in API response:
8- "{property}": "{prop_description}"
9Generate a Python script to verify ’{property}’ in response.
10Rules: {Rules for test gen}
11Format the script as shown: ...’’’’
Figure 6. Constraint Test Generation Prompts
Taking the mined constraints, RBCTest generates test cases for automated verification of API response bodies against their constraints. This process (Figure 7) uses LLMs to convert Request-Response and Response Property constraints into test cases. Each resulting test case is a validation function that takes response data and request parameters as input, and returns whether the response data satisfies their corresponding mined constraints. We also incorporate a semantic verifier that cross-checks test cases against examples in the OAS file. If a test case fails to validate a given correct example, it suggests that the mined constraint may be incorrect.
Figure 7. Constraint Test Generation in RBCTest
Request-Response Constraints Testing
Test cases for Request-Response constraints verify dependencies between a request parameter and its response property. The prompt to generate such test cases (Figure 6) requires four inputs: API input parameter, its mined constraints, corresponding property, and the response schema. To ensure robustness, RBCTest guides the LLM with predefined rules: using a try-catch block for error handling, excluding examples, and following a standardized function template. These rules ensure consistency and focus on constraint verification.
Response Property Constraints Testing
Test cases for Response Property constraints are used to verify response data against their constraints. To create a test case for this type of constraint, RBCTest uses the prompt depicted in Figure 6. This prompt requires three inputs: the response property, its mined constraints, and the response schema. The response property and its mined constraints provide the necessary information for generating validation code, while the response data schema defines the structure and type of the expected data, allowing the LLM to parse response data correctly.
7. Empirical Evaluation
For evaluation, we seek to answer the following questions:
RQ1. [Constraints Mining Accuracy] How well does RBCTest mine the constraints to be used as oracles for verifying API responses compared with the dynamic approach AGORA+ (Alonso2025AGORA_plus)?
RQ2. [Constraints Test Generation Accuracy] How well does RBCTest generate test cases from the mined constraints?
RQ3. [Accuracy in Test Generation from the Correctly Mined Constraints] How accurate is RBCTest in generating test cases for mined constraints?
RQ4. [Usefulness in Response Body Testing] How useful is RBCTest in detecting mismatches between the specification and its working APIs?
RQ5. [Ablation Study] How much do Observation-Confirmation and Semantic Verifier contribute to RBCTest’s performance?
Table 1. Categories of Oracles in RBCTest dataset
Category Test Oracles and Examples
Input-Output Constraints involving the relation between input parameters and output properties. E.g.: input.id_after > return.id
[…]
N-ary atomic Constraints involving two or more variables. E.g.: sellingTotal = total + margins + markup + totalFees - discounts
Composite Constraints involving multiple I/O, single-variable, or atomic constraints. E.g.: input.created_after ¡ output.created_at and output.created_at must isDateTime
AGORA+ dataset (Alonso2025AGORA_plus). This dataset has 7 real-world API services consisting of 11 operations.
RBCTest dataset (rbctest-dataset). We curate the dataset from eight real-world services from GitLab and Stripe, comprising 65 endpoints and 90 operations. These services were selected for several reasons: (1) They feature complex request-response structures across diverse business domains. (2) They have been widely used in API testing research, such as ‘Canada Holidays’ (kat), GitLab-services (yamamoto2021efficient; wu2022combinatorial; atlidakis2019restler; karlsson2020quickrest; lin2022forest; kat), and ‘Stripe’ (kat; martin2022online; mirabella2021deep_stripe). (3) Their active status allows API calls to collect real response data. (4) Their specifications vary in documentation quality, enabling evaluation across different levels of completeness. Stripe, offering a test mode with limited endpoints, contains deeply nested response schemas, often missing values for validation. To mitigate this, we included only Stripe endpoints without nested schemas, retaining 8 in total.
For both datasets, we manually examine the specifications and build the ground truth of correct constraints, i.e., test oracles. During examination, we systematically used the following criteria to look for and derive the correct and sufficient constraints:
(4) Computed Value constraint: The constraint describes how the value of a field is derived from other fields via computation.
(5) Name-based constraint: This primarily focuses on the constraints held within field names like id, code, email, url, uri, etc.
We further referred to the oracles available in the AGORA+ dataset to define and report the oracle categories in the RBCTest dataset (Table 1). The first category is the input/output constraints, which describe the logical relations between the input parameters of APIs and the output properties in their response bodies. The remaining categories contain only the constraints/oracles on the output properties in the response bodies. Specifically, the second category contains the constraints on the individual variables in the output. This category contains unary constraints belonging to several different data types and templates (string, boolean, float, time, url, email, etc). The third category contains the $N$-ary (binary or more) atomic constraints, i.e., involving two or more variables (Table 1). The atomic constraints include comparisons, membership/type checks, and predicate calls without any AND, OR, or universal/existential quantifiers (negations, if any, are pushed inside as part of the atomic). The last category is the super set of the three preceding categories. A composite constraint is expressed as a conjunctive/disjunctive form of the two or more preceding categories, i.e., logical AND and OR operations are applied among these categories.
8. Experimental Results
8.1. Constraint Mining Accuracy (RQ1)
8.1.1. Procedure
For the AGORA/AGORA+ dataset, we ran RBCTest with GPT-4o under the temperature of zero and top_p=0.95 to obtain constraints and manually evaluated them against the ground truth. For the AGORA+ tool, we directly used the results reported in its TOSEM paper (Alonso2025AGORA_plus). For the RBCTest dataset, we ran our tool in the same setting as with the other dataset. For the AGORA+ tool, we checked its publicly available repository and used RESTest (MartinLopez2021Restest) configuration set by AGORA+’s authors to generate 50 request-response pairs following their online instruction. Using this configuration (rather than the default RESTest configuration) ensures the realistic pairs for AGORA+ in RBCTest’s dataset. We chose 50 pairs since its authors reported the best tradeoff between performance and running time at this setting compared to other pair counts.
Table 2 shows samples of constraints identified by RBCTest and invariants generated by AGORA+, illustrating their differences and limited compatibility. To enable meaningful comparison, we adopt a variable-based matching approach. In this context, a variable refers to any request parameter, response property, or intermediate value defined within the API specification. In AGORA+, each invariant is associated with a set of variables, as in Table 2, and a given set of variables may be linked to one or more invariants (e.g. Invariants 3-4). In RBCTest, each constraint refers to one or more variables.
[…]
Table 2. RBCTest’s constraints and AGORA+’s invariants
@@ 8.1.1. Procedure
7 return.ci_forward_deployment_enabled = true
@@ 8 An amount is a positive integer
9 return.milestone = null
10 return.labels[] == []
11 return.statistics.wiki_size one of { 0, 41943 }
8.1.2. Results on the AGORA+ dataset
Table 3. Constraints Mining on the AGORA+ dataset (RQ1)
API RBCTest AGORA+ Overlap Unique
GT TP FP FN P(%) R(%) No.I No. TP P(%) +S +D Eq S D
G.C 100 94.6 7.4 4.4 92.7 95.6 193 97 97 100 6.8 4.0 49.4 34.4 36.8
G.G 157 149 4.2 7.0 97.3 94.5 145 79 68 86.1 2.8 3.0 43.6 99.6 18.6
A.M 59 43.8 4.4 9.2 90.9 77.9 137 63 31 49.2 8.8 2.6 3.6 28.8 16.0
M.C 50 35.4 4 9.2 89.8 74.1 104 55 31 56.4 6.2 1.0 8.0 20.2 15.8
O.I 15 11.8 0.8 2.4 93.7 78.7 17 14 12 85.7 3.2 1.0 2.2 5.4 5.6
O.S 6 5.2 1.8 0.8 74.3 86.7 6 4 3 75.0 1.0 1.0 1.0 2.2 0.0
S.C 27 19.8 6.4 5.4 75.6 74.4 41 21 21 100 6.6 1.0 2.0 10.2 11.4
S.T 21 17.8 4.0 4.2 81.7 80.2 50 24 19 79.2 2.8 5.8 3.8 5.4 6.6
S.A 16 15.4 3.4 1.0 81.9 93.9 68 24 19 79.2 0.0 3.8 8.2 3.4 7.0
Y.B 4 3.2 1.8 0.0 64.0 80.0 60 14 8 57.1 0.0 0.0 1.0 2.2 7.0
Y.V 161 134.2 8.4 23.0 94.1 84.5 132 78 49 62.8 7.2 0.0 10.4 116.6 31.4
Tot. 616 530.2 46.6 65.8 85.1 83.7 953 473 358 75.7 45.4 23.2 133.2 328.4 156.2
  • G.C: Github_createOrgRepo, G.G: Github_getOrgRepo, A.M: AmadeusHotel_getMultiHotel, M.C: Marvel_getComicIndividual, O.I: OMDB_byIdOrTitle, O.S: OMDB_bySearch, S.C: Spotify_createPlaylist, S.T: Spotify_getAlbumTracks, S.A: Spotify_getArtistAlbums, Y.B: Yelp_getBusinesses, Y.V: YouTube_listVideos, GT: # constraints in the ground truth, (# of RBCTest constraints grouped by variables=TP+FP). No.I: #r of invariants. No: # of invariants grouped by variables. +S: RBCTest better, +D: AGORA+ better, Eq: Equivalent. S: RBCTest, D: AGORA+.
As seen in Table 3, RBCTest identified an average of 576.8 constraints out of 616, with an average of 530.2 TPs, resulting in a precision of 85.1% and a recall of 83.7%. AGORA+ detected 953 invariants, and 473 invariants after variable-based grouping. Of these, 358 are TPs, resulting in 75.7% precision.
8.1.3. Results on the RBCTest dataset
Table 4. Constraints Mining on the RBCTest dataset (RQ1)
API RBCTest AGORA+ Overlap Unique
GT TP FP FN P(%) R(%) No.I No. TP P(%) +S +D Eq S D
C.H 42 40.8 6.0 1.2 87.2 97.1 25 23 23 100 3.2 0.0 7.2 30.4 12.6
G.B 61 46.2 3.2 14.8 93.5 75.7 215 123 91 74.0 7.8 6.0 17.8 14.6 59.4
G.C 86 67.0 3.2 19.4 95.4 77.5 313 178 128 71.9 11.2 3.2 25.4 27.2 88.2
G.G 119 78.2 5.6 41.0 93.3 65.6 550 318 225 70.8 17.6 4.0 38.2 18.4 165.2
G.I 268 156.6 3.2 111.8 98.0 58.3 1239 637 451 70.8 38.2 10.0 44.2 64.2 358.6
G.P 248 177.6 4.6 71.2 97.5 71.4 1228 621 376 60.5 30.8 2.0 66.2 78.6 277.0
G.R 55 44.0 3.4 11.0 92.8 80.0 220 122 88 72.1 7.2 2.0 19.2 15.6 59.6
S 107 88.4 18.8 20.8 82.5 81.0 123 86 62 72.1 27.4 1.6 18.4 41.0 14.6
Tot. 986 698.8 48.0 291.0 93.6 70.6 3913 2108 1444 74.0 143.4 28.8 236.6 290.0 1,035.2
  • CH: Canada Holidays, GB: GitLab Branch, GC: GitLab Commit, GG: GitLab Groups, GI: GitLab Issues, GP: GitLab Project, GR: GitLab Repository, S: Stripe.GT: # constraints in the ground truth, (# of RBCTest constraints grouped by variables=TP+FP). No.I: #r of invariants. No: # of invariants grouped by variables. +S: RBCTest better, +D: AGORA+ better, Eq: Equivalent. S: RBCTest, D: AGORA+.
We manually reviewed API specifications and identified 986 correct constraints (Table 4). Our tool identified an average of 746.8 constraints with a precision of 93.6% and a recall of 70.6%. All standard deviations are below 5%.
8.1.4. Overlapping Analysis between RBCTest and AGORA+ results
For the AGORA+ dataset, as seen in Table 3, among overlapping ones, 328.4 constraints were uniquely detected by RBCTest, while 156.2 invariants were uniquely identified by AGORA+. Among the shared ones, on average, 45.4 from RBCTest are better while 23.2 from AGORA+ are better, indicating that RBCTest and AGORA+ are complementary and also share detected constraints.
For the RBCTest dataset, regarding the overlapping results in Table 4, RBCTest and AGORA+ are also complementary, as RBCTest produces 143.4 better oracles with 290.0 uniquely detected ones, and AGORA+ has 28.8 better oracles with 1,035.2 unique ones.
An example of such complementary nature is the Spotify API. Its documentation states that three predefined image sizes (640, 300, 64) are served, but also notes that “the exact size may vary.” Here, AGORA+ detected three explicit values, while RBCTest identified the parameter as an integer, complementing each other’s findings.
8.1.5. Qualitative Analysis and Comparison
[…]
We also examine the cases in which AGORA+ performed well while RBCTest struggled. These cases typically involve invariants that validate URL formats, handle fixed-length tokens, and capture inter-attribute relationships—areas where RBCTest often treats such attributes as plain strings without information. The analysis reveals that RBCTest is limited in these scenarios since it relies primarily on shallow attribute descriptions, typically consisting of the attribute name and a simple type declaration ({type:string}). For example, when an attribute such as return.self is described only with a type, RBCTest is unable to infer underlying constraints.
In contrast, AGORA+ leverages richer execution context and semantic information to detect the kinds of value-specific constraints. These constraints that AGORA+ detected better belong to three main types. Type-based constraints ensure that an attribute conforms to a specific data type, such as verifying that return.owner.gists_url or return.self is of type Url. Value-based constraints impose conditions on numeric or string values, for instance, enforcing that LENGTH(return.runners_token) == 29 or that return.stats.deletions >= 0 and does not exceed return.stats.total. Finally, inter-attribute constraints capture relationships among multiple fields, including equality (return.owner.gists_url == return.organization.gists_url) and substring relationships (e.g., return.owner.url being a substring of both return.owner.followers_url and return.owner.following_url).
Table 5. Types of Constraints detected by RBCTest
No. Category/Test Oracles Count  No. Category/Test Oracles Count
1 I/O 309   9 isBoolean 18
2 isUrl 290   10 isNumber 18
3 isDateTime 160   11 isUnixTime 14
4 Value-In-Set 150   12 Template Literals 11
5 Composite 87   13 ArrayTypeOracle_isString 3
6 isDate 45   14 Array_Specific_Sizes 3
7 String_Specific_Length 35   15 N-ary atomic 3
8 Value-In-Range 29   16 isTime 2
8.1.6. Analysis on Types of Constraints Detected by RBCTest
As shown in Table 5, we summarize the distribution of 1,177 true positive constraints identified across 16 main categories in both the AGORA and RBCTest datasets. Among these, the I/O constraint type, which represents logical and dependency relationships between input and output fields within individual APIs, accounts for 309 instances (26.3%). This indicates that RBCTest is effective at mining commonly described I/O relationships within APIs.
The isUrl category constitutes the second largest group, comprising 290 constraints (24.6%), showing the widespread presence of URL-related attributes in the API specifications. The isDateTime constraints, which specify date/time formats and rules, rank next with 160 constraints (13.6%). The Value-In-Set category, containing 150 constraints (12.7%), reflects the common use of enumerated or bounded value restrictions that encode domain-specific semantics such as status codes, country abbreviations, or numeric thresholds in the APIs. Following these, the Composite category includes 87 constraints (7.4%), showing that RBCTest can capture complex logical conditions composed of multiple atomic constraints.
The remaining categories account for about 15.4% of all detected constraints, covering specialized rules such as date formats, string lengths, numeric ranges, and array checks — reflecting finer-grained or domain-specific properties. This result demonstrates that RBCTest can detect a wide variety of common oracles found in real-world API schemas and response bodies.
8.2. Constraint Test Generation Accuracy (RQ2)